# FineMath: A Fine-Grained Mathematical Evaluation Benchmark for Chinese Large Language Models

### v1

@@ Abstract
| Yan Liu1, Renren Jin1, Lin Shi2, Zheng Yao3, Deyi Xiong1††thanks: Corresponding author. |
@@ 5.4. Results across 17 MWP Categories
| Question |
学校买来120本课外读物,其中的40%分给高年级,剩下的课外读物按5:3分给中年级和低年级,低年级分到多少本? |
The school bought 120 supplementary books, 40% of which were given to the upper grades. The remaining books were divided between the middle and early grades in a ratio of 5:3. How many books did the early grades receive? |
| Prompt 0+Question |
@@
GPT-4 Response: The number of supplementary books allocated to the upper grades is 120*40%=48 books. The remaining books total 120-48=72 books. The supplementary books allocated to the early grades are calculated as 72*3/(5+3)=27 books. Answer: The early grades are allocated 27 books. |
| Prompt 1+Question |
@@
GPT-4 Response: 36 books. Firstly, the school allocates 40% of the 120 books to the upper grades. This totals to 120*40%=48 books. The remaining books amount to 120-48 = 72 books. Then, these 72 books are divided between the middle and early grades at a ratio of 5:3. The total of this ratio is 5+3=8. Therefore, the number of books allocated to the early grades is 72*(3/8) =27 books. |
| Prompt 4+Question |
@@
GPT-4 Response: 36 books |
| Prompt 0+Question |
GPT-4 Response: C:27 |
GPT-4 Response: C:27 |
| (with MCQs) | ||
| A:72 B:48 C:27 D:80 |
@@
| Prompt 0+Question |
GPT-4 Response: D:36 |
GPT-4 Response: D:36 |
| (with D replaced) | ||
| A:72 B:48 C:27 D:36 |
@@ 6.1. Prompts Really Does Matter
Prompt like “Answer:” appear to encourage the model to forego reasoning and directly provide the answer, which increases the likelihood of generating incorrect responses. An example is shown in Table 8: Prompt 4+Question.
@@
Comparison results are displayed in Table 9. We can observe a significant difference in accuracy between option prediction (with multiple-choice questions) and direct answer generation, with a gap that can exceed 10%. Interestingly, restructuring the task in the form of multiple-choice questions seems to reduce the accuracy of high-performing models while increasing the accuracy of models that perform poorly. Upon examining instances, we have found that the answer option can act as another type of prompt influencing the model’s performance. For example in Table 8: Prompt 0+Question (with MCQs) and Prompt 0+Question (with D replaced).
Examples of GPT-4 outputs with different prompts and task forms are shown in Table 8. We want to understand why GPT-4 would provide the incorrect answer “36” under Prompt 1. Therefore, we utilize Prompt 3 to have GPT-4 explain its reasoning for choosing “36”. Interestingly, GPT-4 mentions the correct number “27” in the explanation, but still provides the incorrect answer, “36”. Given the seeming importance of “36”, we replace one of the options in the multiple-choice question with “36”. The model, which initially selected the correct answer, abandoned it in favor of “36”. Further examples can be observed in the responses generated by other models. For instance, the reasoning process may be incorrect, yet the correct answer is ultimately chosen. The options provided to the model also seem to influence the model’s generation probability to a certain degree. Therefore, we recommend using generation, rather than option prediction, for a more accurate evaluation of LLMs.
@@ 8. Bibliographical References
[…]

### v2

@@ Abstract
Yan Liu1, Renren Jin1, Ling Shi2, Zheng Yao3, Deyi Xiong1††thanks: Corresponding author. |
@@ 5.4. Results across 17 MWP Categories
Question | 学校买来120本课外读物,其中的40%分给高年级,剩下的课外读物按5:3分给中年级和低年级,低年级分到多少本? | The school bought 120 supplementary books, 40% of which were given to the upper grades. The remaining books were divided between the middle and early grades in a ratio of 5:3. How many books did the early grades receive? | |
@@
Prompt 0+Question | GPT-4 Response: C:27 | GPT-4 Response: C:27 | |
(with MCQs) | |||
@@
Prompt 0+Question | GPT-4 Response: D:36 | GPT-4 Response: D:36 | |
(with D replaced) | |||
@@ 6.1. Prompts Really Does Matter
Prompt like “Answer:” appear to encourage the model to forego reasoning and directly provide the answer, which increases the likelihood of generating incorrect responses. An example is shown in Table 5.4: Prompt 4+Question.
@@
Comparison results are displayed in Table 9. We can observe a significant difference in accuracy between option prediction (with multiple-choice questions) and direct answer generation, with a gap that can exceed 10%. Interestingly, restructuring the task in the form of multiple-choice questions seems to reduce the accuracy of high-performing models while increasing the accuracy of models that perform poorly. Upon examining instances, we have found that the answer option can act as another type of prompt influencing the model’s performance. For example in Table 5.4: Prompt 0+Question (with MCQs) and Prompt 0+Question (with D replaced).
Examples of GPT-4 outputs with different prompts and task forms are shown in Table 5.4. We want to understand why GPT-4 would provide the incorrect answer “36” under Prompt 1. Therefore, we utilize Prompt 3 to have GPT-4 explain its reasoning for choosing “36”. Interestingly, GPT-4 mentions the correct number “27” in the explanation, but still provides the incorrect answer, “36”. Given the seeming importance of “36”, we replace one of the options in the multiple-choice question with “36”. The model, which initially selected the correct answer, abandoned it in favor of “36”. Further examples can be observed in the responses generated by other models. For instance, the reasoning process may be incorrect, yet the correct answer is ultimately chosen. The options provided to the model also seem to influence the model’s generation probability to a certain degree. Therefore, we recommend using generation, rather than option prediction, for a more accurate evaluation of LLMs.
@@ 8. Bibliographical References
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners. In Advances in Neural Information Processing Systems 33: Annual Conference on Neural Information Processing Systems 2020, NeurIPS 2020, December 6-12, 2020, virtual.